OpenStudy (anonymous):

What is the value of 6^2 × 6^3? A. 216 B. 7,776 <--- C. 46,656 D. 248,832 @hero Can you check this one too? :) Can you also give me all the formulas? :)

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

\(6^{2}\times 6^{3}=6^{5}=7776\)

OpenStudy (anonymous):

According to PEMDAS, exponents first. \[6^2=6\times6=36\] \[6^3=6\times6\times6=216\] And multiply. \[36\times216=7,776\] So, you are correct.

OpenStudy (anonymous):

PEMDAS: Parenthesis Exponents Multiplication & Division Addition & Subtraction *Huge Note: For MD and AS, do whatever comes from left to right. So, in: \[6\div6\times1\] the division part would come first.